OpenStudy (nevillesnetwork):

Medal & Fan Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received: x − 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive? Do not enter comma for placeholder values.

OpenStudy (steve816):

Hey there! Since y is the total number of votes cast, all you have to do is plug in 125,000 for y.

OpenStudy (steve816):

Basically it would be x - 62,500 = 0.25(125,000) Then, solve for x!

OpenStudy (nevillesnetwork):

OH! Thanks so much!

OpenStudy (nevillesnetwork):

Can I try to solve this and then you check it? @steve816

OpenStudy (steve816):

No problem! And sure, I'll check your answer.

OpenStudy (nevillesnetwork):

Yay! Ok, give me a moment.

OpenStudy (nevillesnetwork):

x-62,500 = 31250 x-62,500 -31250 = 31250-31250 I got x - 31250 but I don't think thats correct... I missed something, or did I not? @steve816

OpenStudy (nevillesnetwork):

@steve816 Plz check this, I only have a few minutes... :O

OpenStudy (steve816):

x - 62,500 = 0.25(125,000) x - 62,500 = 31250 Now from here, you have to add 62,500 to both sides. x = 93750